[figure: 9.7 cm]

THERMODYNAMICS OF THE SPIN-1/2 HEISENBERG CHAIN

WITH NEXT-NEAREST NEIGHBOR INTERACTIONS

S. Dyeyev Institute of Physics, Polish Academy of Sciences Al. Lotnikow 32/46 PL-02-668 Warsaw, Poland

Linear chains of N spin-\frac{1}{2} with nearest and next-nearest exchange interactions have been investigated by numerical diagonalization of the full energy matrix. The diagonalization has been performed for ferro- and antiferromagnetic coupling between next-nearest as well as nearest neighbors. The magnetic susceptibility and specific heat versus temperature have been calculated. The behavior of the thermal and magnetic properties of finite chain has been estimated by extrapolation (N ⟶ ∞). Various physical properties such as magnetic susceptibility and specific heat were compared with experimental data for Pb[Cu(SO_{4})(OH)_{2}] natural single crystal.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : S. Dyeyev

Address for correspondence : Sergiy Dyeyev Institute of Physics, Polish Academy of Sciences Al. Lotnikow 32/46 PL-02-668 Warsaw, Poland

Email address : sergd@ifpan.edu.pl